If Masterston, Inc. just paid $2.55 to its shareholders as the annual dividend. Simultaneously, the company announced that future dividends will be increasing by 5.5 percent per year, indefinitely and investors require a rate of return of 11 percent on this stock. What is the current price? What is the price in three years? What is the price in fifteen years?
Answer #1

Current price=D1/(Required return-Growth rate)

=(2.55*1.055)/(0.11-0.055)

=$48.91(Approx).

P3=Current price*(1+Growth Rate)^3

=48.91*(1.055)^3

=$57.44(Approx).

P15=Current price*(1+Growth Rate)^15

=48.91*(1.055)^15

=$109.20(Approx).